Let D be an integral domain. Then prove that there exists a 

field F that contains a subring isomorphic to D.


1
Expert's answer
2021-02-24T06:09:38-0500

Solution:

Proof:

Let a, b ∈ D with b"\ne" 0.

Since E is a field, "b^{-1}" ∈ E.

Thus define the function φ : F → E as φ(a, b) = "ab^{\u22121}" .

Here we show that φ is well defined, so suppose (a, b),(a'b') ∈ F such that (a, b) = (a'b'). (Note that b, b'"\ne" 0.)

Then by the definition of equivalence classes in the field of quotients, we have ab' = ba' ⇒ "ab^{\u22121}" = "a'b'^{\u22121}" ⇒ φ((a, b)) = φ((a'b' )).

Hence φ is well-defined.

Next, we show that φ preserves addition and multiplication.

Let (a, b),(c, d) ∈ F (b, d '"\ne"0).

Then φ((a, b)+(c, d)) = φ((ad+bc, bd)) = (ad+bc)("d^{-1}b^{-1}" ) = "add^{\u22121} b^{-1}+bcd^{\u22121} b^{\u22121} = ab^{\u22121}+cd^{\u22121}" = φ((a, b))+φ((c, d)) and φ((a, b)(c, d)) = φ((ac, bd)) = "acd^{\u22121} b ^{\u22121} = ab^{\u22121} cd^{\u22121}" = φ((a, b))φ((c, d))

Thus φ preserves the operations and is a homomorphism.

Note that φ : F → φ(F) is onto by definition.

Next, we show that φ is one-to-one. Suppose that φ((a, b)) = φ((c, d)). Then "ab^{\u22121} = cd^{\u22121}" which implies ad = bc.

Hence (a, b) = (c, d) by the definition of equivalence classes in the field of quotients and therefore φ is one-to-one. Since φ is a bijection, φ is an isomorphism.

We know that φ(F) is a field by its construction.

φ(F) is an integral domain (D) plus the inverses of the elements of D. Therefore we have a subfield of E, specifically φ(F) that is isomorphic to F.

Hence, proved.